40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1047     Entry time: Tue Oct 14 19:18:18 2008
Author: Yoichi 
Type: Update 
Category: Computers 
Subject: BootFest 
Rana, Yoichi

Most of the FE computers failed around the lunch time.
We power cycled those machines and now all of them are up and running.
I confirmed that the both arms lock.
Now the IFO is in "Restore last auto-alignment" status.
ELOG V3.1.3-